# Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1

### v3

@@ 2.5 Statistics
ChEBI-20 (Edwards et al., 2021) Targeted Generation 26,407 3,000
PubChem (Liu et al., 2023) Targeted Generation 12,000 2,000
@@
Table 3: Comparison with existing text-based molecule generation benchmarks and datasets. Notably, we denote the statistics of the description guided molecule design part in Mol-Instructions.
@@
3 Experiments
3.1 Models
@@
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024b), Claude-3 (Anthropic, 2024a), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General LLMs. This group contains open-source LLMs that are tuned with the instruction following capability, which can be used for a wide range of tasks and applications. Specifically, we benchmark Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), and Gemma3-12B (Team et al., 2025). These LLMs are top-tier open-source general LLMs within our computation budget.
@@
OpenMolIns Fine-tuned LLMs. We further fine-tune LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama3.1-8B-Instruct on OpenMolIns dataset for comparison. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M because the model has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small size model can also help us study the data scaling law within reasonable budget. Furthermore, Llama3.2-1B-Instruct and Llama3.1-8B-Instruct are also selected due to their advanced capabilities and similar architectures.
Figure 2: The performance of LLMs benchmarked in S2-Bench. LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models of unknown parameters are denoted as horizontal lines.
3.2 Findings
The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations. Our finding suggests that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery. Open-source general LLMs can often lack chemical corpora during their pre-training stage. As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results. The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding. In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning. Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rates of 39.33%. More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved a weighted average success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct. This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
Figure 3: Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama-3.1-8B-Instruct.
F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing. As shown in Figure 3, our experiments with Galactica-125M across five data scales of OpenMolIns provide new insights into the data scaling law for molecule generation. Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, indicating that tasks involving complex, de novo synthesis strongly benefit from larger and more diverse training data. By contrast, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
Interestingly, when examining Llama-3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks. This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
4 Conclusion
In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation. By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery. S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization. Our comprehensive benchmarking of 28 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
Appendix A Related Work
Molecule discovery is a cornerstone of scientific progress, underpinning advances in both drug development and material design (Du et al., 2022). The integration of artificial intelligence into this process has driven a paradigm shift in the pharmaceutical landscape, substantially improving the efficiency and accuracy of identifying and developing novel therapeutic candidates.
Early work in this direction employed transformer-based models such as MolT5 (Edwards et al., 2022), which leveraged large-scale self-supervised learning to produce high-quality SMILES strings from textual prompts. Building on this foundation, models like KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023) integrated molecular graphs and biochemical texts to jointly enhance molecular understanding and generation. More recently, 3D-MoLM (Li et al., 2024d) advanced the field by incorporating spatial configurations, thereby improving the geometric validity and accuracy of generated molecular structures.
Parallel to these architecture-driven improvements, LLMs such as MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) have demonstrated strong in-context learning capabilities, enabling adaptive molecule generation by retrieving and utilizing relevant examples from the provided context. Besides, MolReFlect (Li et al., 2024c), emphasizes fine-grained alignment between molecular structures and natural language descriptions through a teacher–student paradigm, which effectively captures subtle structural–semantic relationships. Moreover, Mol-R1 (Li et al., 2025) takes a step further by introducing DeepSeek-R1 like Long-CoT reasoning in molecule generation, enabling robust and interpretable natural language-driven molecule generation.
Appendix B Implementation Details
We implement various scripts to facilitate the testing of the aforementioned models. For proprietary models, we adopt the OpenAI API [^1] https://platform.openai.com/docs/ framework. For open-source general LLMs, we utilize both the VLLM [^2] https://docs.vllm.ai framework and the OpenAI framework. For the remaining LLMs, we adopt the Hugging Face transformers library [^3] https://huggingface.co/docs/ for inference.
Furthermore, BioT5 is designed to use SELFIES, an alternative string representation of molecules by using special tokens to ensure valid generations (Krenn et al., 2019), as input instead of SMILES. Consequently, we convert the molecule SMILES strings into SELFIES format on BioT5.
Appendix C Hyper Parameters
@@
Table 4: Hyper-parameters.
Appendix D Data Construction
In this section, we introduce the construction details of S2-Bench and OpenMolIns dataset, as well as the prompt templates. Figure 4 demonstrates the overall data construction process for S2-Bench as well as the OpenMolIns. Generally, we utilize Zinc250K to construct the test samples in S2-Bench, while adopting PubChem for the training samples in OpenMolIns. Notably, when constructing the OpenMolIns, we exclude all the molecules in Zinc250K to avoid data leakage and ensure the novelty score of the generated molecules.
Figure 4: Data construction workflow of S2-Bench & OpenMolIns.
D.1 MolEdit
For the MolEdit task, we consider the common operations on modifying functional groups in a given molecule (i.e., add, drop, and substitute), which are simple tasks for human experts but challenging to LLMs. In this case, we further develop three corresponding subtasks: AddComponent, DelComponent, and SubComponent. Prompt templates for MolEdit are shown in Table 5. However, there are different kinds of functional groups, and some functional groups can play an important role in the molecule structure, such as connecting two separate parts of the molecule, which makes them unsuitable for these operations above as these operations will entirely change the structure of the molecule. In this case, we aim to make a slight change in the molecule structure and limit most of the functional groups we choose within the end groups.
Prompt Templates for MolEdit
AddComponent
Please add a {} to the molecule {}.
Modify the molecule {} by adding a {}.
Add a {} to the molecule {}.
DelComponent
Please remove a {} from the molecule {}.
Modify the molecule {} by removing a {}.
Remove a {} from the molecule {}.
SubComponent
Please substitute a {} in the molecule {} by {}.
Modify the molecule {} by replacing a {} by {}.
Replace a {} in the molecule {} by {}.
Please replace a {} in the molecule {} with {}.
Modify the molecule {} by substituting a {} with {}.
Substitute a {} in the molecule {} with {}.
Table 5: Prompt Templates for MolEdit.
Table 6 presents the functional groups that are taken into account for AddComponent and DelComponent, along with their respective selection weights. Given that certain functional groups are encountered less frequently in real-world scenarios, we have implemented a weighted random selection process for AddComponent, which ensures that less common functional groups are assigned a lower probability to be chosen.
Functional Group benzene ring hydroxyl aldehyde carboxyl amide
Weights 15 15 5 5 10
Functional Group amine nitro halo nitrile thiol
Weights 5 5 5 1 1
Table 6: Functional Groups that are considered in AddComponent and DelComponent, as well as their weights to be selected in AddComponent.
For SubComponent, our focus is exclusively on end groups for simplification, which include hydroxyl, aldehyde, carboxyl, nitro, halo, nitrile, and thiol, which ensures that the editing operations are confined to substituting the existing functional group with another from this list, thereby maintaining the integrity of the molecule’s overall structure without altering it fundamentally.
D.2 MolOpt
MolOpt is designed to optimize molecular properties through the refinement of molecule structures, is not a brand-new task. Previously, GNN-based methods have been widely adopted in this task, while these methods can only help with one specific subtask at a time. In contrast, S2-Bench requires one single LLM to optimize molecules with different metrics and optimization directions. In this work, we specifically focus on enhancing specific characteristics that are crucial for drug discovery and chemical synthesis, including LogP, MR, and QED. The prompt templates for MolOpt are listed in Table 7.
Prompt Templates for MolOpt
LogP
Please optimize the molecule {} to have a lower/higher LogP value.
Modify the molecule {} to decrease/increase its LogP value.
Optimize the molecule {} to have a lower/higher LogP value.
Please modify the molecule {} to decrease/increase its LogP value.
Modify the molecule {} to have a lower/higher LogP value.
MR
Please optimize the molecule {} to have a lower/higher MR value.
Modify the molecule {} to decrease/increase its MR value.
Optimize the molecule {} to have a lower/higher MR value.
Please modify the molecule {} to decrease/increase its MR value.
Modify the molecule {} to have a lower/higher MR value.
QED
Please optimize the molecule {} to have a lower/higher QED value.
Modify the molecule {} to decrease/increase its QED value.
Optimize the molecule {} to have a lower/higher QED value.
Please modify the molecule {} to decrease/increase its QED value.
Modify the molecule {} to have a lower/higher QED value.
Table 7: Prompt Templates for MolOpt.
Below are the detailed explanations of the chemical properties adopted in the three subtasks of MolOpt:
LogP refers to the logarithm of the partition coefficient, which is a measure of a molecule’s hydrophilicity or lipophilicity. It is an important factor in determining a compound’s bioavailability and membrane permeability.
Molecular Refractivity (MR) is a measure of the molar refractive index, which provides insight into the molecular size and the degree of molecular branching. It is used to assess the overall shape and bulk of a molecule.
Quantitative Estimation of Drug-Likeness (QED) is a computational metric that evaluates the drug-likeness of a molecule based on a set of predefined rules. A higher QED score suggests a greater likelihood that the molecule will have favourable pharmacological properties.
Among the three subtasks, the QED optimization, which assesses the potential of a molecule to become a drug for curing diseases, is generally considered the most challenging property for molecule optimization. In contrast, LogP and MR can be more directly inferred from the molecule’s structure, thus making them more straightforward to optimize.
D.3 MolCustom
To enable customized design of molecules, we consider three fundamental features for describing the molecule, including atoms, bonds, and functional groups. Given the specified category and number of atoms, bonds, and functional groups, LLMs should generate the molecule as we request. The prompt templates for MolCustom are shown in Table 8. Below, we present the construction details of the three subtasks for MolCustom:
Prompt Templates for MolCustom
AtomNum
Please generate a molecule with {} atom(s).
Please generate a molecule composed of {} atom(s).
Please generate a molecule consisting {} atom(s).
The molecule has {} atom(s).
The molecule is composed of {} atom(s).
The molecule consists of {} atom(s).
There is a molecule with {} atom(s).
There is a molecule composed of {} atom(s).
There is a molecule consisting of {} atom(s).
The molecule contains {} atom(s).
BondNum
Please generate a molecule with {} bond(s).
Please generate a molecule composed of {} bond(s).
Please generate a molecule consisting {} bond(s).
The molecule has {} bond(s).
The molecule is composed of {} bond(s).
The molecule consists of {} bond(s).
There is a molecule with {} bond(s).
There is a molecule composed of {} bond(s).
There is a molecule consisting of {} bond(s).
The molecule contains {} bond(s).
FunctionalGroup
Please generate a molecule with {} group(s).
Please generate a molecule composed of {} group(s).
Please generate a molecule consisting {} group(s).
The molecule has {} group(s).
The molecule is composed of {} group(s).
The molecule consists of {} group(s).
There is a molecule with {} group(s).
There is a molecule composed of {} group(s).
There is a molecule consisting of {} group(s).
The molecule contains {} group(s).
Table 8: Prompt Templates for MolCustom.
Atom carbon oxygen nitrogen sulfur fluorine chlorine bromine iodine phosphorus
Weights [Mandatory] 5 3 3 2 2 2 2 1
Atom boron silicon selenium tellurium arsenic antimony bismuth polonium
Weights 1 1 1 1 1 1 1 1
Table 9: Atoms that are considered in AtomNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal atoms in the molecule structures.
AtomNum. Table 9 shows the atoms we consider in AtomNum, as well as their weights to be selected. Notably, carbon, as the basic unit in organic chemicals, is a mandatory option. The number of carbon atoms ranges from 1 to 40, while the number of other selected atoms ranges from 1 to 5. This setting relieves the difficulty for generation, as LLMs could generate a carbon backbone first and attach the remaining atoms to the backbone one by one.
Bond single double triple rotatable aromatic
Weights 5 4 3 1 1
Table 10: Chemical bonds that are considered in BondNum, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal bonds in the molecule structures.
BondNum. Similarly, we select five different kinds of chemical bonds: single, double, triple, rotatable, and aromatic, as shown in Table 10. For the single bond, if selected, the number can vary from 1 to 50. For the aromatic bond, the number follows the rules of the formation of aromatic bonds, varying from 5 to 20. Moreover, the number of these remaining bonds, if selected, is specified from 1 to 5.
Functional Group benzene ring hydroxyl anhydride aldehyde ketone carboxyl ester amide amine nitro
Weights 15 15 2 5 5 10 5 5 5 2
Functional Group halo thioether nitrile thiol sulfide disulfide sulfoxide sulfone borane
Weights 2 1 1 1 1 1 1 1 1
Table 11: Functional Groups that are considered in FunctionalGroup, as well as their weights to be selected. Higher weights indicate that they are more likely to be selected. For simplification, we only consider normal and edge functional groups in the molecule structures.
FunctionalGroup. Lastly, we also specify functional groups in the molecule structure. Table 11 shows the range of functional groups and their weights that are taken into consideration.
Notably, in MolCustom, if not required, LLMs can generate any number of these unspecified atoms, bonds, and functional groups. However, for these specified items, LLMs should strictly follow the requirements.
Appendix E Testing Process

### v4

@@ 3.5. Statistics
Table 3. Comparison with existing text-based molecule generation benchmarks and datasets. Notably, we denote the statistics of the description-guided molecule design part in Mol-Instructions (Fang et al., 2024).
@@
ChEBI-20 (Edwards et al., 2022) Targeted Generation 26,407 3,000
PubChem324K (Liu et al., 2023) Targeted Generation 12,000 2,000
@@
We also report molecular distribution statistics for Zinc-250K, PubChem, and our pre-selected samples used in MolEdit and MolOpt. Specifically, we compare the average (i) atom count, (ii) ring count, (iii) branch count, and (iv) path length across datasets.
Table 4. Molecular Distribution Statistics.
Dataset # Samples Avg Atom Count Avg Ring Count Avg Branch Count Avg Path Length
Zinc-250K 250,000 23.15 2.76 7.31 12.48
PubChem 9,000,000 25.18 2.8 7.87 13.07
S²-Bench’s sampling 30,000 23.17 2.74 7.34 12.47
As shown in Table 4, Zinc-250K exhibits broadly similar statistical trends to PubChem, suggesting that Zinc-250K captures the core characteristics of drug-like chemistry and remains a reasonable proxy for discovery-scale small molecules. Meanwhile, the distribution of our pre-selected samples aligns closely with the original Zinc-250K, suggesting that our pre-selection does not substantially distort the underlying molecular distribution.
4. Experiments
4.1. Models
@@
Proprietary Models. This category includes LLMs that are only accessible via commercial API services. In this work, we benchmark GPT-4o, GPT-4-turbo, GPT-3.5-turbo (OpenAI, 2023), Claude-3.5 (Anthropic, 2024a), Claude-3 (Anthropic, 2024b), and Gemini-1.5-pro (Deepmind, 2024), which are all the most advanced LLMs with powerful reasoning and generalization capabilities. These LLMs are pre-trained on vast pre-training corpora, normally including chemical-related knowledge.
Open-source General & Chemical LLMs. This group contains open-source general LLMs with instruction following capability for a wide range of tasks, including Llama3-70B-Instruct, Llama3-8B-Instruct, Llama3.1-8B-Instruct, Llama3.2-1B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-R1-distill-Qwen-7B (Guo et al., 2025a), Qwen2-7B-Instruct (Yang et al., 2024), yi-1.5-9B (Young et al., 2024), chatglm-9B (GLM et al., 2024), Gemma3-12B (Team et al., 2025). Additionally, domain-specific LLMs like ChemLLM-20B (Zhang et al., 2024), SmileyLlama (Cavanagh et al., 2026),and ChemDFM-v1.5-8B (Zhao et al., 2025) are also considered.
@@
OpenMolIns Fine-tuned LLMs. We further fine-tune LLMs like Galactica-125M (Taylor et al., 2022), Llama3.2-1B-Instruct, and Llama3.1-8B-Instruct on OpenMolIns dataset for comparison. We specifically include the experiments on five distinct data sizes of OpenMolIns for Galactica-125M because the model has been pre-trained on scientific corpora and has proved its effectiveness in molecule-related tasks (Liu et al., 2023). Meanwhile, a small-sized model can also help us study the data scaling law within a reasonable budget. Furthermore, Llama3.2-1B-Instruct and Llama3.1-8B-Instruct are also selected due to their advanced capabilities and similar architectures.
4.2. Hyperparameters
@@
Figure 3. The performance of LLMs evaluated in S2-Bench. LLMs fall into 4 categories: Proprietary Models, Open-source General LLMs, Open-source ChEBI-20 Fine-tuned LLMs, and OpenMolIns Fine-tuned LLMs. Models with unknown parameters are denoted as horizontal lines.
4.3. Findings
The precise control, in fact, is crucial for molecular discovery, as many properties of interest, such as solubility, toxicity, and active site composition, depend on specific structural configurations. Our findings suggest that open-domain natural language–driven molecule generation task can serve as an effective diagnostic tool, revealing systematic weaknesses in current LLMs and providing rich failure cases that can inform future instruction tuning or reinforcement learning strategies to improve molecule structural constraint adherence.
F2: While pre-training provides a foundation for LLM performance, instruction tuning is indispensable for guiding and optimizing their capabilities in molecular discovery. Open-source general LLMs can often lack chemical corpora during their pre-training stage. As a result, models like Qwen2-7B-Instruct and yi-1.5-9B obtained unsatisfactory results. The observed shortcoming can be attributed to the lack of chemical data in pre-training, despite their notable proficiency in mathematics and general language understanding. Surprisingly, open-source chemical LLMs such as ChemLLM-20B and ChemDFM-v1.5-8B achieve relatively poor performance, indicating that current chemical LLM pre-training still fails to capture the molecular structural details. In contrast, Llama3.1-8B-Instruct delivers superior performance at a relatively modest model scale, demonstrating a more balanced and comprehensive capability across diverse domains.
Meanwhile, our experiments with OpenMolIns highlight the effectiveness of instruction tuning. Fine-tuned on the largest OpenMolIns scale (i.e., xlarge), Llama3.1-8B-Instruct surpassed all other LLMs with an average weighted success rate of 39.33%. More remarkably, the much smaller model, Galactica-125M, within just 125 million parameters, achieved an average weighted success rate of 25.73%, outperforming models two orders of magnitude larger, such as Llama3-70B-Instruct. This remarkable outcome indicates that task-specific instruction tuning on a large, high-quality corpus can be highly effective, suggesting a practical avenue for LLMs to efficiently narrow the performance gap in molecular discovery.
Table 6. Comparing WSR and SR with Human Expert Evaluation across BioT5-base and Galactica-125M. WSR consistently achieves stronger alignment with expert judgment than SR alone. Spearman correlations are computed over the six subtasks per model; $^{*}p<0.05$.
BioT5-base Galactica-125M Human Expert Eval.
Task Subtask SR WSR SR WSR BioT5-base Galactica-125M
AddComponent 0.3462 0.0542 0.5842 0.3423 0.64 3.57
MolEdit DelComponent 0.1668 0.0266 0.6526 0.3318 0.22 3.00
SubComponent 0.0684 0.0108 0.1872 0.1128 0.19 2.60
LogP 0.5158 0.0787 0.7362 0.4429 0.58 3.74
MolOpt MR 0.5060 0.0808 0.7124 0.4059 0.67 3.47
QED 0.5068 0.0801 0.5786 0.3285 0.64 3.40
Spearman $\rho$ (vs. Human Expert Eval.) 0.551 $\mathbf{0.899}^{*}$ 0.714 $\mathbf{0.886}^{*}$
p-value 0.257 0.015 0.111 0.019
Figure 4. Task-specific performance scaling with increasing data in S2-Bench. (a) Galactica-125M. (b) Llama3.1-8B-Instruct.
F4: The data scaling law does not always hold: task-specific bottlenecks limit gains in molecule optimization and editing. As shown in Figure 4, our experiments with Galactica-125M across the five data scales provide new insights into the data scaling law for molecule generation. Overall, we observe that larger datasets generally improve performance, but the effect is highly task-dependent.
In MolCustom, scaling the dataset from large to xlarge yields a remarkable 286% average performance gain, while there is limited performance improvement from light to large, indicating that tasks involving complex, de novo synthesis strongly benefits from larger and more diverse training data. Conversely, Galactica-125M obtains only modest gains in MolOpt, and even negative gains in MolEdit, suggesting that additional data alone cannot overcome existing bottlenecks in these relatively simpler tasks.
Interestingly, when examining Llama3.1-8B-Instruct, we find that performance continues to increase when scaling from the large to xlarge data level, except for LogP and MR subtasks. This contrast generally highlights that small models like Galactica-125M are capacity-limited, whereas larger models can more effectively leverage additional data.
4.4. Validating WSR with Human Evaluation
We conducted an extra correlation analysis to justify our design for using similarity in Weighted Success Rate (WSR). Here, we focused on the unusual performance of BioT5-base in MolEdit and MolOpt, where the model achieves high success rates but performs poorly in the quality term, indicating that the generated molecules are not modified from the original molecule but from scratch.
For this analysis, we first randomly sampled 100 examples from each subtask and then asked human experts to evaluate the correctness and quality of the generated molecules. Each example is assigned an expert score from 0 to 5, where:
  • 5 denotes a perfect success,
  • 4 means success but with redundant operations (a little different structure),
  • 3 shows failure but modified from the original molecule (similar structure),
  • 2 denotes failure and a little different structure,
  • 1 indicates success but an unrelated/different structure,
  • 0 denotes failure and an unrelated/different structure.
We calculate the average score of expert evaluation. Table 6 presents per-subtask SR, WSR, and expert scores for BioT5-base and Galactica-125M, with Spearman correlations reported at the bottom. WSR achieves significantly stronger alignment with expert judgment than SR alone across both models: for BioT5-base, $\rho_{\mathrm{WSR}}=0.899$ ($p=0.015$) versus $\rho_{\mathrm{SR}}=0.551$ ($p=0.257$); for Galactica-125M, $\rho_{\mathrm{WSR}}=0.886$ ($p=0.019$) versus $\rho_{\mathrm{SR}}=0.714$ ($p=0.111$). In both cases WSR reaches statistical significance while SR does not.
Therefore, it is essential to apply similarities as weights of Success Rate to reflect the genuine capabilities of LLMs more faithfully.
The results confirm that SR is a noisy proxy for quality: as a binary criterion, it rewards any molecule satisfying the target constraint irrespective of structural distance from the input, allowing models to score high SR by generating chemically valid but structurally unrelated outputs, a pattern human experts consistently penalize. WSR corrects for this by down-weighting successes achieved through large structural departures, and the significantly higher Spearman correlations confirm that this re-weighting tracks human judgment more faithfully.
4.5. Extending MolOpt to Additional Properties
The choice of LogP, MR, and QED in MolOpt was motivated by the need for automated, reproducible, and unambiguous evaluation at scale, which currently rules out many complex properties. To address the concern that this set may be overly simplistic, we extend MolOpt with two more challenging properties: Topological Polar Surface Area (TPSA), widely used to estimate drug permeability, and HOMO-LUMO Gap (HLG), a fundamental quantum-chemical descriptor of molecular reactivity. Due to resource constraints, we report results on a set of representative models in Table 7. The overall success rates on TPSA and HLG are noticeably lower than those on LogP/MR/QED, confirming that these properties pose a harder optimization challenge and that MolOpt is far from saturated; meanwhile, no single model dominates across both properties, suggesting complementary strengths on different property axes and reinforcing the value of MolOpt as a discriminative benchmark that naturally extends beyond the original three properties.
Table 7. Extended MolOpt results on TPSA and HOMO-LUMO Gap (HLG). SR denotes Success Rate and WSR denotes Weighted Success Rate.
Model TPSA SR TPSA WSR HLG SR HLG WSR
MolT5-large 0.3568 0.0347 0.3460 0.0319
ChemDFM-v1.5-8B 0.0634 0.0440 0.0920 0.0578
Llama3.1-8B-Instruct 0.3654 0.1745 0.1760 0.0881
5. Conclusion
In this study, we introduced S2-Bench and OpenMolIns, the first benchmark and instruction tuning dataset for evaluating the capabilities of LLMs in open-domain natural language-driven molecule generation. By moving beyond traditional one-to-one text-to-molecule mappings, our benchmark accommodates the open-ended nature of molecular discovery. S2-Bench focuses on the realistic molecular reasoning and design, rather than simple pattern matching and memorization. Our comprehensive benchmarking of 31 LLMs not only highlights the challenges faced by current models and the limitations of existing targeted molecule generation datasets, but also demonstrates the substantial potential of LLMs for natural language–driven molecule discovery.
Acknowledgments
This research was supported by the National Natural Science Foundation of China (Grant No.: T2541073, 62372314), the Hong Kong Research Grants Council under the Theme-based Research Scheme (project no. T41-517/25-N), and the NSFC/RGC Joint Research Scheme (N_PolyU5179/25). The experimental part of this work was supported by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
Appendix A Testing Process